# Entanglement entropy of two disjoint intervals and spin structures in interacting chains in and out of equilibrium

### 15 Dec 2023 v1

@@ 6.1 Tensor network implementation
In this section we describe the tensor network algorithms that we used to obtain numerical data for the tripartite information of three adjacent subsystems in the spin chains of interest. Let |$\ket{\psi}$ by the ground state of a given Hamiltonian $H$. After fixing a local basis of the Hilbert space associated to each spin, we can expand the state as follows
@@
Once the variational ansatz (6.8) is chosen, the next step is to develop an algorithm to find the parameters that minimize the energy. We use the variational uniform matrix product states (VUMPS) algorithm, formulated in Ref. VUMPSfirst to perform variational energy minimization in infinite 1D quantum systems. Ref. VUMPSfirst shows that the VUMPS algorithm generally outperforms its more-famous predecessors, the infinite Denisy Matrix Renormalization Group (iDMRG) iDMRG and the infinite Time Evolving Block Decimation (iTEBD) iTEBD_first , both in convergence speed and accuracy of the approximation. It also seems to be a good choice for studying the critical systems we are interested in. The algorithm comes with two main schemes of convergence to the variational minimum: sequential and parallel. The original paper VUMPSfirst showed that these two approaches are typically equivalent in terms of performance and our simulations are based on the sequential algorithm. In practice, we use the Julia-language implementation of the algorithm available in the work-in-progress package ITensorInfiniteMPS.jl, based on the ITensors.jl library ITensor . All the results that we show in this work have been obtained setting the bond dimension to $\chi=140$.
@@
6.2 Numerical analysis
@@
Figure 10: Contributions of different spin structures $(\varepsilon,\delta)$ to the Rényi-$2$ tripartite information of three finite adjacent subsystems $A,B,C$ in the XXZ spin chain on an infinite chain. Function $\mathcal{F}_{f}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ (see Eq. (3.18) and Eq. (3.19)) is shown versus the cross ratio $x=|A||C|/[(|A|+|B|)(|B|+|C|)]$ for all four torus spin structures. The lengths $|A|=|C|$ are varied, while the length of $|B|$ is set to $6$ (filled markers) or $10$ (open markers). The points are obtained using the method of fusion of models, discussed in the main text, incorporating the data from the tensor network simulations. The solid lines display the prediction (2.10) for the limit of large subsystem lengths.
We have used the described procedure with the interleaved model (6.2) to compute the functions $\mathcal{F}_{f}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ in the ground state of the XXZ spin chain (2.1), for all spin structures $\varepsilon,\delta$ of the torus, for different configurations $(A,B,C)$. We have chosen the parameters $(\Delta,h)=(-0.3,0)$ of the XXZ spin chain for all our simulations. In Fig. 4 we show the results for configurations with fixed cross ratio (1.4), for two different values, $x=1/4$ and $x=1/2$. In Fig. 10 we show the results for configurations with the length of the central subsystem $B$ fixed to two different values. In both figures the results are compared with the prediction (2.10) for $\alpha=2$. Our numerical results agree very well with the predictions, providing the first numerical check of its kind in an interacting model. This also provides an indirect check of the Rényi-$2$ tripartite information in the ground state of the fermionic XXZ chain (5.1) as well as in the ground state of the Jordan-Wigner-$\nu$ deformations (2.3) for any $\nu$.
@@ 7 Insights into the tripartite information after global quenches
One of the commonest protocols of global quench starts with preparing the system in the ground state of a local, translationally invariant Hamiltonian, which is then suddenly perturbed, for example, by changing a global magnetic field. In generic situations the system is expected to locally thermalise, in the sense that the expectation values of local observables will relax to thermal values. An exception to that behaviour, which has attracted much interest, is when the system has infinitely many (pseudo)local conservation laws, which generally means that the system is integrable. The stationary state can then be characterised by a so-called generalised Gibbs ensemble (GGE) Rigol2007Relaxation ; Doyon2017Thermalization ; Vidmar2016Generalized ; Essler2016Quench . A GGE shares many properties in common with thermal states, for example, subsystems are generally described by reduced density matrices with an extensive entropy. It can however also exhibit exceptional features. Specifically, static correlations in thermal states of systems with local interactions decay exponentially with the distances; in a generalised Gibbs ensemble of a system with local interactions, that is not always the case. Ref. Maric2022Universality considered some settings in which one should expect also algebraically decaying correlations at late times after a global quench in an integrable model. It was observed observed that such exceptional property leaves clear marks in the entanglement entropies of subsystems. In particular, Refs Maric2022Universality ; Maric2023Universality ; Maric2023universality2 studied the tripartite information after global quenches in noninteracting spin chains and identified contributions that are generally killed by the quench and others that are instead only mildly affected by nonequilibrium time evolution. As a result, the tripartite information of large adjacent subsystems is not trivialised by the limit of infinite time and, remarkably, remains a function of the cross ratio such as in conformal critical ground states. It was also pointed out that the final result has a universal property: when the tripartite information survives the quench, the limit $x\rightarrow 1^{-}$ turns out to be nonzero and, in the large class of noninteracting models investigated in Ref. Maric2023Universality , it approaches a universal $-\log 2$.
@@ 8 Conclusions
The results for the fermionic XXZ chain generalise the works of Refs Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement on the entanglement entropy of disjoint blocks in non-interacting hopping chains, which are described by free massless Dirac fermions. While for the latter the tripartite information is zero for any $\alpha$, the tripartite information in the fermionic XXZ chain, described by the massless Thirring QFT with NS boundary conditions on all cycles, exhibits a more interesting behavior. In particular, it can be both positive and negative.
@@
There are several open problems that our work brings to light, particularly in the direction of quantum quenches. First, our conjecture about the asymptotic behaviour of the tripartite information after a small quench in an interacting integrable model should be confirmed and generalised to larger quenches by an exact calculation of time evolution after the quench. Second, Ref. Hayden2013Holographic argued that a negative tripartite information should suggest a domination of quantum entanglement over classical correlations. Since we confirmed the negativity of the residual tripartite information after global quenches, it is natural to wonder whether that is measuring a quantum or classical feature (or both of them). To that aim it might be useful to study some other quantity in which the distinction between classical and quantum correlations is sharper, for example the entanglement negativity Vidal2002 ; Calabrese2012negativityPRL . We also wonder whether the residual tripartite information might be connected to some unusual property of the entanglement Hamiltonian Dalmonte2022 of disjoint subsystems.
@@ Appendix C Noninteracting systems: free fermion techniques
Such a simplicity is related to the property that the reduced density matrix (3.2) is a fermionic Gaussian, i.e. an operator of the form $\rho(\Gamma)\propto\exp{\left(\sum_{j,\ell}a_{j}^{\dagger}W_{j,l}a_{\ell}/4\right)}$, where $W$ is an antisymmetric (in general complex) matrix. Clearly, the moments appearing in Eq. (3.19) can be then computed using $\mathrm{tr}\rho_{A}^{\alpha}=\exp[(1-\alpha)S_{\alpha}(A)]$.
@@ Appendix D About the effective parameters in the interleaved model
Figure 11: The same as in Fig. 4a, where open point markers correspond to data obtained using the parameters $\gamma,h$ of the interleaved model (6.2) in Eq. (6.4), whereas filled point markers correspond to renormalised data. The latter data are obtained by choosing effective parameters in the XY subchain so as to match the numerical expectation values of two local observables. The R-R is not affected by the change, but the other spin structures are very sensitive to it and only the renormalised data show good agreement with the predictions. See the discussion around Eq. (6.5) and Sec. 6.2.
@@ Appendix E From the tripartite information of Jordan-Wigner spin deformations to XXZ spin structure
$e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}=\sum_{n=1}^{M_{\alpha}}g_{\alpha,n}\gamma_{\alpha,n}^{\nu}-\sum_{n=1}^{\tilde{M}_{\alpha}}\tilde{g}_{\alpha,n}(-\gamma_{\alpha,n})^{\nu}\,,$
where each $\gamma$ is a ratio of partition functions and $g_{\alpha,n}$ are positive integers that represent the degeneracy of the terms. We can generally express $\{\gamma_{\alpha,n}\}$ in terms of $\{e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}\}$ if we consider a large enough number of deformations. If $g_{\alpha,n}=1$ and $\tilde{g}_{\alpha,n}=0$, we can readily express $\{\gamma_{\alpha,n}\}$ as the solutions to the polynomial
@@
$e^{\tilde{I}_{3}^{(2;j)}(x)}=e^{I_{3}^{(2;j)}(x)}+(-1)^{j}\tfrac{\gamma_{2,4}^{j}(x)}{2}$ (E.2)
@@
$e_{1}=$ $2e^{\tilde{I}_{3}^{(2;1)}(x)}=2e^{I_{3}^{(2;1)}(x)}-\gamma_{2,4}(x)$ (E.3)
$e_{2}=$ $2e^{2\tilde{I}_{3}^{(2;1)}(x)}-e^{\tilde{I}_{3}^{(2;2)}(x)}=2e^{2I_{3}^{(2;1)}(x)}-e^{I_{3}^{(2;2)}(x)}-2e^{I_{3}^{(2;1)}(x)}\gamma_{2,4}(x)$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-(2e^{2I_{3}^{(2;1)}(x)}-e^{I_{3}^{(2;2)}(x)})\gamma_{2,4}(x)\,.$
@@
$\gamma_{3,1}^{j}+3\gamma_{3,2}^{j}+3\gamma_{3,3}^{j}+3\gamma_{3,4}^{j}-6(-1)^{j}\gamma_{3,5}^{j}=4e^{2I_{3}^{(3;j)}(x)}\,,$ (E.10)
@@
$\gamma_{3,1}^{j}+\sum_{k=1}^{3}(\gamma_{3,2_{k}}^{j}+\gamma_{3,3_{k}}^{j}+\gamma_{3,4_{k}}^{j})=4e^{2I_{3}^{(3,j)}(x)}+\sum_{k=1}^{6}(-\gamma_{3,5_{k}})^{j}\,,$ (E.11)
@@
$e_{n}[\{\gamma_{3,1},\gamma_{3,2_{\bullet}},\gamma_{3,3_{\bullet}},\gamma_{3,4_{\bullet}}\}]=\\ \frac{4^{n}}{n!}\det\left|\begin{matrix}e^{2I_{3}^{(3;1)}}-\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\tfrac{1}{4}&0&\dots&\\ e^{2I_{3}^{(3;2)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{2}}{4}&e^{2I_{3}^{(3;1)}}-\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\frac{2}{4}&0\dots\\ \vdots&&\ddots&\ddots\\ e^{2I_{3}^{(3;n-1)}}+\tfrac{\sum_{k}(-\gamma_{3,5_{k}})^{n-1}}{4}&e^{2I_{3}^{(3;n-2)}}+\tfrac{\sum_{k}(-\gamma_{3,5_{k}})^{n-2}}{4}&\dots&e^{2I_{3}^{(3;1)}}-\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\frac{n-1}{4}\\ e^{2I_{3}^{(3;n)}}+\tfrac{\sum_{k}(-\gamma_{3,5_{k}})^{n}}{4}&e^{2I_{3}^{(3;n-1)}}+\tfrac{\sum_{k}(-\gamma_{3,5_{k}})^{n-1}}{4}&\dots&e^{2I_{3}^{(3;2)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{2}}{4}&e^{2I_{3}^{(3;1)}}-\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}\end{matrix}\right|\,.$ (E.12)
@@
$4e^{2I_{3}^{(3;10+m)}}+\sum_{k=1}^{6}(-\gamma_{3,5_{k}})^{10+m}+\\ \sum_{n=1}^{10}(-1)^{n}e_{n}[\{\gamma_{3,1},\gamma_{3,2_{\bullet}},\gamma_{3,3_{\bullet}},\gamma_{3,4_{\bullet}}\}](4e^{2I_{3}^{(3;10+m-n)}}+\sum_{k=1}^{6}(-\gamma_{3,5_{k}})^{10+m-n})=0\,.$ (E.13)

### 11 Mar 2024 v3

@@ 6.1 Tensor network implementation
In this section we describe the tensor network algorithms that we used to obtain numerical data for the tripartite information of three adjacent subsystems in the spin chains of interest. Let |$\ket{\psi}$ be the ground state of a given Hamiltonian $H$. After fixing a local basis of the Hilbert space associated to each spin, we can expand the state as follows
@@
Once the variational ansatz (6.7) is chosen, the next step is to develop an algorithm to find the parameters that minimize the energy. We use the variational uniform matrix product states (VUMPS) algorithm, formulated in Ref. VUMPSfirst to perform variational energy minimization in infinite 1D quantum systems. Ref. VUMPSfirst shows that the VUMPS algorithm generally outperforms its more-famous predecessors, the infinite Density Matrix Renormalization Group (iDMRG) iDMRG and the infinite Time Evolving Block Decimation (iTEBD) iTEBD_first , both in convergence speed and accuracy of the approximation. It also seems to be a good choice for studying the critical systems we are interested in. The algorithm comes with two main schemes of convergence to the variational minimum: sequential and parallel. The original paper VUMPSfirst showed that these two approaches are typically equivalent in terms of performance and our simulations are based on the sequential algorithm. In practice, we use the Julia-language implementation of the algorithm available in the work-in-progress package ITensorInfiniteMPS.jl, based on the ITensors.jl library ITensor . The results that we show in this work have been obtained setting the bond dimension up to $\chi=180$.
@@
6.2 On the numerical data
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(0.1,0)$. The three subsystems have equal lengths $|A|=|B|=|C|$, so the cross ratio is $x=1/4$. The dashed lines show the free case ($\Delta=0$) for comparison.
(a) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(0.1,0)$. The three subsystems have equal lengths $|A|=|B|=|C|$, so the cross ratio is $x=1/4$. The dashed lines show the free case ($\Delta=0$) for comparison.
(b) Contribution from different spin structures for $\alpha=2$ and $(\Delta,h)=(0.1,0)$. The subsystem lengths are set to be $(|A|,|B|,|C|)=(2\ell,\ell,3\ell)$ for $\ell=1,2,3,\ldots$, so the cross ratio is $x=1/2$. The lines R-NS and NS-R overlap. The dashed lines show the free case for comparison.
Figure 10: The same as in Fig. 4, but for $(\Delta,h)=(0.1,0)$. The factor $(-1)^{4\varepsilon\delta}$ is now omitted since the prediction for the R-R term is positive.
@@ 6.2 On the numerical data
Figure 11: Contributions of different spin structures $(\varepsilon,\delta)$ to the Rényi-$2$ tripartite information of three finite adjacent subsystems $A,B,C$ in the infinite XXZ spin chain with $\Delta=-0.3$. Function $\mathcal{F}_{f}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ (see Eq. (3.18) and Eq. (3.19)) is shown versus the cross ratio $x=|A||C|/[(|A|+|B|)(|B|+|C|)]$ for all four torus spin structures. The lengths $|A|=|C|$ are varied, while the length of $|B|$ is set to $6$ (filled markers) or $10$ (open markers). The points are obtained using the method of fusion of models, incorporating the data from the tensor network simulations. The solid lines are the predictions (2.10), whereas dashed lines are the corresponding contributions in the free case ($\Delta=0$).
Figure 12: The Rényi-$2$ tripartite information in the infinite fermionic XXZ chain (2) with $\Delta=-0.3$ for finite adjacent subsystems $A,B,C$. The lengths $|A|=|C|$ are varied, while the length of $|B|$ is set to $6$ (filled markers) or $10$ (open markers). The points are obtained using the method of fusion of models, incorporating the data from the tensor network simulations. The solid line is the prediction (2.4), whereas the dashed line is the Rényi-$2$ tripartite information in the fermionic XX chain ($\Delta=0$), which is zero for any $x$.
We have used the described procedure with the interleaved model (6.1) to compute the functions $\mathcal{F}_{f}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ in the ground state of the XXZ spin chain (2.1) for all spin structures $\varepsilon,\delta$ of the torus, for different configurations $(A,B,C)$ and different values of $\Delta$. We discuss here the data for $(\Delta,h)=(-0.3,0)$ and $(\Delta,h)=(0.1,0)$, for which we reached a satisfactory numerical convergence and are exemplar of the general behaviour observed.
Figures 4 and 10 show the data at fixed cross ratio (1.4) equal to either $x=1/4$ or $x=1/2$. In Fig. 11, instead, the length of the central subsystem $B$ is fixed to two different values and the other lengths are set to the same value, varying over a range that is numerically accessible with high enough accuracy. Finally, Fig. 12 shows a comparison between prediction and numerical data for the fermionic Rényi-$2$ tripartite information.
In all the cases, the data are in good visual agreement with the predictions (2.10). Note that the larger disagreement for smaller $x$ in Fig. 11 and Fig. 12 should be expected, in that, for fixed $|B|$, a smaller cross ratio corresponds to smaller $|A|=|C|$ and, therefore, to larger finite-size effects. Despite the overall good visual agreement, we point out that, in order to check the agreement quantitatively, one would need to work out the asymptotics of the finite-size corrections, which could depend on the particular spin structure and would not necessarily match the leading corrections expected in the tripartite information. We have not addressed this issue, therefore we leave a real quantitative check to future investigations. Even if our analysis is more qualitative than quantitative, it provides anyway the first numerical check of its kind in an interacting model. It also provides an indirect check of the Rényi-$2$ tripartite information in the ground state of the Jordan-Wigner-$\nu$ deformations (2.3) for any $\nu$.
@@ 7 Insights into the tripartite information after global quenches
One of the commonest protocols of global quench starts with preparing the system in the ground state of a local, translationally invariant Hamiltonian, which is then suddenly perturbed, for example, by changing a global magnetic field. In generic situations the system is expected to locally thermalise, in the sense that the expectation values of local observables will relax to thermal values. An exception to that behaviour, which has attracted much interest, is when the system has infinitely many (pseudo)local conservation laws, which generally means that the system is integrable. The stationary state can then be characterised by a so-called generalised Gibbs ensemble (GGE) Rigol2007Relaxation ; Doyon2017Thermalization ; Vidmar2016Generalized ; Essler2016Quench . A GGE shares many properties in common with thermal states, for example, subsystems are generally described by reduced density matrices with an extensive entropy. It can however also exhibit exceptional features. Specifically, static correlations in thermal states of systems with local interactions decay exponentially with the distances Bluhm2022exponentialdecayof ; in a generalised Gibbs ensemble of a system with local interactions, that is not always the case. Ref. Maric2022Universality considered some settings in which one should expect also algebraically decaying correlations at late times after a global quench in an integrable model. It was observed that such exceptional property leaves clear marks in the entanglement entropies of subsystems. In particular, Refs Maric2022Universality ; Maric2023Universality ; Maric2023universality2 studied the tripartite information after global quenches in noninteracting spin chains and identified contributions that are generally killed by the quench and others that are instead only mildly affected by nonequilibrium time evolution. As a result, the tripartite information of large adjacent subsystems is not trivialised by the limit of infinite time and, remarkably, remains a function of the cross ratio such as in conformal critical ground states. It was also pointed out that the final result has a universal property: when the tripartite information survives the quench, the limit $x\rightarrow 1^{-}$ turns out to be nonzero and, in the large class of noninteracting models investigated in Ref. Maric2023Universality , it approaches a universal $-\log 2$.
@@ 8 Conclusions
The results for the fermionic XXZ chain generalise the works of Refs Casini2005 ; Casini2009reduced_density ; Casini2009Entanglement on the entanglement entropy of disjoint blocks in non-interacting hopping chains, which are described by the free massless Dirac fermion. While for the latter the tripartite information is zero for any $\alpha$, the tripartite information in the fermionic XXZ chain, described by the massless Thirring QFT with NS boundary conditions on all cycles, exhibits a more interesting behavior. In particular, it can be both positive and negative.
@@
There are several open problems that our work brings to light, particularly in the direction of quantum quenches. First, our conjecture about the asymptotic behaviour of the tripartite information after a small quench in an interacting integrable model should be confirmed and generalised to larger quenches by an exact calculation of time evolution after the quench. Second, Ref. Hayden2013Holographic argued that a negative tripartite information should suggest a domination of quantum entanglement over classical correlations. Since we have confirmed that the residual tripartite information after global quenches is negative, it is natural to wonder whether that is measuring a quantum or classical feature (or both of them). To that aim it might be useful to study some other quantity in which the distinction between classical and quantum correlations is sharper, for example the entanglement negativity Vidal2002 ; Plenio2005 ; Calabrese2012negativityPRL . We also wonder whether the residual tripartite information might be connected to some unusual property of the entanglement Hamiltonian Dalmonte2022 of disjoint subsystems.
@@ Appendix C Noninteracting systems: free fermion techniques
Such a simplicity is related to the property that the reduced density matrix (3.2) is a fermionic Gaussian, i.e. an operator of the form $\rho(\Gamma)\propto\exp{\left(\sum_{j,\ell}a_{j}^{\dagger}W_{j,l}a_{\ell}/4\right)}$, where $W$ is an antisymmetric matrix. Clearly, the moments appearing in Eq. (3.19) can be then computed using $\mathrm{tr}\rho_{A}^{\alpha}=\exp[(1-\alpha)S_{\alpha}(A)]$.
@@ Appendix D About the effective parameters in the interleaved model
Figure 13: The same as in Fig. 4a, where open point markers correspond to data obtained using the parameters $\gamma,h$ of the interleaved model (6.1) in Eq. (6.3), whereas filled point markers correspond to renormalised data. The latter data are obtained by choosing effective parameters in the XY subchain so as to match the numerical expectation values of two local observables. The R-R is not affected by the change, but the other spin structures are very sensitive to it and only the renormalised data show good agreement with the predictions. See the discussion around Eq. (6.4).
@@ Appendix E From the tripartite information of Jordan-Wigner spin deformations to XXZ spin structure
$e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}=\sum_{n=1}^{M_{\alpha}}g_{\alpha,n}\gamma_{\alpha,n}^{\nu}-\sum_{n=M_{\alpha}+1}^{M_{\alpha}+\tilde{M}_{\alpha}}\tilde{g}_{\alpha,n}\gamma_{\alpha,n}^{\nu}\,,$ (E.1)
where each $\gamma_{\alpha,n}$ denotes the contribution of a different spin structure $\mathcal{F}_{f}^{(\mathrm{XXZ})}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)$, $M_{\alpha}$ and $\tilde{M}_{\alpha}$ are integers denoting the limits of the sum and $g_{\alpha,n}$ and $\tilde{g}_{\alpha,n}$ are positive integers that represent the degeneracy of the terms. We can generally express $\{\gamma_{\alpha,n}\}$ in terms of $\{e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}\}$ if we consider a large enough number of deformations. If $g_{\alpha,n}=1$ and $\tilde{g}_{\alpha,n}=0$, we can readily express $\{\gamma_{\alpha,n}\}$ as the solutions to the polynomial
@@
$e^{\tilde{I}_{3}^{(2;j)}(x)}=e^{I_{3}^{(2;j)}(x)}+\tfrac{\gamma_{2,4}^{j}(x)}{2}$ (E.3)
@@
$e_{1}=$ $2e^{\tilde{I}_{3}^{(2;1)}(x)}=2e^{I_{3}^{(2;1)}(x)}+\gamma_{2,4}(x)$ (E.4)
$e_{2}=$ $2e^{2\tilde{I}_{3}^{(2;1)}(x)}-e^{\tilde{I}_{3}^{(2;2)}(x)}=2e^{2I_{3}^{(2;1)}(x)}-e^{I_{3}^{(2;2)}(x)}+2e^{I_{3}^{(2;1)}(x)}\gamma_{2,4}(x)$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(2e^{2I_{3}^{(2;1)}(x)}-e^{I_{3}^{(2;2)}(x)})\gamma_{2,4}(x)\,.$
@@
$\gamma_{3,1}^{j}+3\gamma_{3,2}^{j}+3\gamma_{3,3}^{j}+3\gamma_{3,4}^{j}-6\gamma_{3,5}^{j}=4e^{2I_{3}^{(3;j)}(x)}\,,$ (E.11)
@@
$\gamma_{3,1}^{j}+\sum_{k=1}^{3}(\gamma_{3,2_{k}}^{j}+\gamma_{3,3_{k}}^{j}+\gamma_{3,4_{k}}^{j})=4e^{2I_{3}^{(3,j)}(x)}+\sum_{k=1}^{6}\gamma_{3,5_{k}}^{j}\,,$ (E.12)
@@
$e_{n}[\{\gamma_{3,1},\gamma_{3,2_{\bullet}},\gamma_{3,3_{\bullet}},\gamma_{3,4_{\bullet}}\}]=\\ \frac{4^{n}}{n!}\det\left|\begin{matrix}e^{2I_{3}^{(3;1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\tfrac{1}{4}&0&\dots&\\ e^{2I_{3}^{(3;2)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{2}}{4}&e^{2I_{3}^{(3;1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\frac{2}{4}&0\dots\\ \vdots&&\ddots&\ddots\\ e^{2I_{3}^{(3;n-1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{n-1}}{4}&e^{2I_{3}^{(3;n-2)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{n-2}}{4}&\dots&e^{2I_{3}^{(3;1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}&\frac{n-1}{4}\\ e^{2I_{3}^{(3;n)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{n}}{4}&e^{2I_{3}^{(3;n-1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{n-1}}{4}&\dots&e^{2I_{3}^{(3;2)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}^{2}}{4}&e^{2I_{3}^{(3;1)}}+\tfrac{\sum_{k}\gamma_{3,5_{k}}}{4}\end{matrix}\right|\,.$ (E.13)
@@
$4e^{2I_{3}^{(3;10+m)}}+\sum_{k=1}^{6}\gamma_{3,5_{k}}^{10+m}+\\ \sum_{n=1}^{10}(-1)^{n}e_{n}[\{\gamma_{3,1},\gamma_{3,2_{\bullet}},\gamma_{3,3_{\bullet}},\gamma_{3,4_{\bullet}}\}](4e^{2I_{3}^{(3;10+m-n)}}+\sum_{k=1}^{6}\gamma_{3,5_{k}}^{10+m-n})=0\,.$ (E.14)